# Specific heat of Gd3+ and Eu2+-based magnetic compounds

### v3

@@ Abstract
We have studied theoretically the specific heat of a large number of non-frustrated magnetic structures described by the Heisenberg model for systems with total angular momentum $J=7/2$, corresponding to the 4f7 configuration of Gd+3 and Eu+2. For a given critical temperature (determined by the magnitude of the exchange interactions), we find that, to a high degree of accuracy, the specific heat is governed by two primary parameters: the effective number of neighbors, which dictates the extent of spatial and quantum fluctuations, and the axial anisotropy. Using these two parameters we fit the specific heat of four Gd compounds and two Eu compounds, achieving a remarkable agreement. Our work provides a framework to describe the specific heat of 4f7 systems within a unified framework.
@@ I Introduction
However, an exception is expected for compounds containing Gd+3 and Eu+2. Both of these ions correspond to the 4f7 configuration. In this case, the ground-state multiplet, constructed by Hund’s rules, is $^{8}S_{7/2}$. This corresponds to a total spin $S=7/2$, total orbital angular momentum $L=0$, and total angular momentum $J=7/2$. We show in this manuscript that the specific heat of 4f7 systems can be characterized primarily by two parameters: the effective number of neighbors $z$ and axial anisotropy $K$. This effective number of neighbors $z$ accounts for the number of interaction bonds of each ion. This simple framework could enable more accurate predictions of thermal properties in magnetic materials and serve as guide for specific model Hamiltonians.
@@
In this work, we report the specific heat of a large number of structures described by a Heisenberg Hamiltonian $H$ containing exchange interactions at different distances and an axial anisotropy term $H_{A}$:
@@
$\mathbf{S}_{i},S_{iz}$ denote the spin (or angular momentum) $|S|=7/2$ operators at site $i$ and its projection, $\delta$ labels the non-equivalent space vectors connecting different spins at short distances (nearest and possibly further neighbors), and $J_{\delta}$ is the corresponding exchange interaction. Couplings $J_{\delta}$ are ferromagnetic if negative or antiferromagnetic if positive. We use $S$ instead of $J$ to denote the operators to avoid confusion with the exchange interactions.
The theoretically studied magnetic structures are non-frustrated and correspond to either ferromagnetic interactions or non-frustrated antiferromagnetic ones. We find that the specific heat of all these structures can be characterized to a large degree of accuracy by two parameters: i) the effective number of neighbors (to be defined more precisely below) and ii) the magnetic anisotropy defined by $H_{A}$. Using these two parameters, and the known value of $T_{C}$, we fit the data for the specific heat of GdNi3Ga9, GdPbBi, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. The corresponding results are presented in Section VI.
I.1 The Effective Number of Neighbors
@@
I.2 Axial anisotropy in 4f7 ions
@@
Nevertheless, due to spin-orbit coupling, the ground-state multiplet acquires a small admixture of the $^{6}P_{7/2}$ state with $L=1$. An explicit calculation for Gd+3 in Appendix B of Ref. 18 estimates that the excited multiplet contributes approximately $3\%$ to the ground state. This admixture introduces a component with $L=1$ in the ground state. It allows the ground state to split under the influence of crystal fields. This occurs if the symmetry at the rare-earth site is not high. Specifically, if there is only a single axis of higher order, the anisotropy term $H_{A}=C[3L_{z}^{2}-L(L+1)]$ becomes relevant.
@@ II Methods
We have calculated the specific heat of a large number of magnetic structures. The approach used a two-point numerical differentiation to compute the derivative of the internal energy as a function of temperature. The internal energy was calculated using quantum Monte Carlo (QMC) simulations with the default “minbounce” algorithm (implemented in the “dirloop_sse” application in the ALPS libraries [19, 20]). QMC allow us to handle large system sizes and complex lattices while properly accounting for spatial and quantum fluctuations that mean-field approaches miss. Simulations were performed on systems of up to $8\times 30^{3}=216000$ magnetic moments. A typical simulation required approximately 10000 CPU hours per lattice on a standard workstation.
@@
To generate the lattice configurations for the ALPS simulations, we employed the “SUNNY” package [21] in conjunction with crystallographic information files (CIF) obtained from public repositories. Specifically, CIF files for Fe in simple cubic (SC), body-centered cubic (BCC), hexagonal close-packed (HCP), and face-centered cubic (FCC) structures, as well as C in diamond, MgAgAs (for GdPbBi), Th(CrSi)2 (for GdCu2Ge2), ErNiSi3 (for GdNiSi3), Mg2MnGa3 (for Eu2Pd3Sn) and MgAl2Cu (for EuPdSn2) were retrieved from the Materials Project [22] database. The ErNi3Al9 CIF file was obtained from the Crystallography Open Database [23] using data from reference 24. Lattice parameters for GdCu2Ge2 were taken from [25], for GdPbBi from [26], for GdNiSi3 from [27], for Eu2Pd3Sn from [28], and for EuPdSn2 from [29]. In each case, the magnetic structure was constructed by assigning a spin moment $S=7/2$ to the respective atoms (C, Fe, Mg, Th, or Er). For all lattices, except HCP, we used a rectangular cuboid supercell geometry as the conventional unit cell, resulting in varying numbers of sites within the supercell, with the diamond lattice requiring the largest (8 sites). Simulations were performed on systems containing $N^{3}$ supercells with periodic boundary conditions, typically with $N=10$, 2020, and $30$. For the critical temperature, results were extrapolated to the thermodynamic limit assuming a $1/N$ dependence.
@@
The well-known sign problem [30] associated with QMC simulations limited our investigation to non-frustrated magnetic structures. At low temperatures, the number of Monte Carlo steps required for statistically reliable results increases exponentially. Consequently, the internal energy was primarily computed in a temperature range between $0.2T_{C}$ and $2T_{C}$, where $T_{C}$ is the critical temperature of the system. In a few cases, simulations were extended to $0.1T_{C}$ at low temperatures. However, the reliability of the results below $0.2T_{C}$ is limited due to finite-size effects. We used $2\times 10^{5}$ Monte Carlo steps for thermal equilibration and $4\times 10^{6}$ steps for measurements. These parameters results in a well converged energy and a smooth specific heat. The resulting error bars (not shown in figures) are close to the line width or fluctuations seen in figure 2.
@@
III Role of the effective number of neighbors
@@
Figure 1: Critical temperature $T_{C}$ as a function of effective number of neighbors $z$ for several ferromagnetic structures. Finite-size results ($N=10,20,30$) and extrapolated values are shown, with the mean-field value $T_{C}^{\mathrm{MF}}$ indicated by the straight line.
@@
In this section we report our results for the specific heat and critical temperature for several simple lattices as a function of the number of neighbors that interact with a given site. For simplicity, in all lattices considered, all sites are equivalent, and we assume all $J_{\delta}=J_{1}$, where $J_{1}$ is the nearest-neighbor (NN) interaction, up to a certain distance $d$ and 0 for longer distances. Therefore, the coordination number given by equation 4 reduces to $\Sigma_{\delta}z_{\delta}. zz$ is given by the total number of neighbors within a sphere of radius $d$. In Section V we discuss more general cases demonstrating that the results are not severely affected by this assumption.
@@
Most of our studies assume ferromagnetic interactions ($J_{\delta}<0$). Generally, the results for the corresponding non-frustrated antiferromagnetic structures are very similar, particularly for $z>6$, except at very low temperatures, where the specific heat, dominated by magnons, scales as $T^{3/2}$ ($T^{3}$) for ferro- and antiferromagnetic structures respectively.
@@
In Fig. 1, we show the critical temperature $T_{C}$ as a function of $z$ for several ferromagnetic structures, including finite-size results and an extrapolated value. Calculations have been done in finite systems containing $N^{3}$ conventional unit cells, with $N=10$, 20, and 30, as explained in Section II. The extrapolated value is also indicated in the figure.
@@
As expected, $T_{C}$ increases almost linearly with $z$ as the mean-field value $T_{C}^{\mathrm{MF}}$ (the effective magnetic field due to the interactions is proportional to $zJ_{1}$, see Section IV.1). But it is lower than $T_{C}^{\mathrm{MF}}$, particularly for low $z$ due to the effect of spatial and quantum fluctuations, which is larger for low $z$.
@@
For large enough $z$, one expects that $T_{C}$ approaches $T_{C}^{\mathrm{MF}}$. However, in finite systems with periodic boundary conditions, the coordination number decreases when the interaction range $d$ becomes comparable to the system size, because neighbors at different distances overlap. This might prevent us from reaching the thermodynamic limit with our method. Nevertheless, arguing as in the introduction that fluctuations decrease as $1/z$ , we should expect that the critical temperature approaches the MF result much faster than the results shown on Fig. 1.
@@
Figure 2: Specific heat $C/T$ versus reduced temperature $T/T_{C}$ for several ferromagnetic structures with different effective neighbor counts $z$. Top panel ($z\leq 12$) and bottom panel ($z\geq 12$) show distinct ranges. The curves for SC, FCC, and HCP structures coincide at $z=18$, while SC, BCC, and EuPdSn2 coincide at $z=26$, explaining the single curves shown for each set. LHc(4) denotes the layered honeycomb lattice with $z=4$. The bottom inset shows finite-size effects for HCP(12) with $N=10$ (black), $N=20$ (red), and $N=30$ (green) in the range $0.5<T/T_{C}<1.5$.
@@
In Fig. 2 we show the results for the specific heat $C$ as a function of temperature for a large number of ferromagnetic structures with different numbers of magnetic neighbors $z$ (the number of nearest neighbors up to third order is shown in Table 1). Also shown is the mean-field result which corresponds to the limit $z\rightarrow\infty$. This result shows larger $C$ for $T<T_{C}$, while $C=0$ for $T>T_{C}$. In contrast, the specific heat for a honeycomb layered system with $z=4$ (three nearest neighbors within the plane of the honeycomb layer and one above or below the layer in alternating sites, LHc(4)) shows the smallest specific heat below $T_{C}$ and the largest one above $T_{C}$. Note that the diamond structure, which is distinct from the previous one but also has $z=4$, has a very similar specific heat. The low-temperature shoulder is a Schottky-like effect due to the internal field [32, 14]. The lower panel inset shows that finite size effects are only appreciable near $T_{C}$ as exemplified in the case of the HCP(12) lattice. For this reason, we show specific heat curves from QMC corresponding to $N=20$.
@@
Increasing $z$ from 4 to 6, the specific heat of the simple cubic structure with six NN displays the expected trend of increasing (decreasing) $C$ below (above) $T_{C}$. Increasing $z$ up to 14 using the BCC structure with 6 NN and 8 next-nearest neighbors (NNN), the same trend is observed in general. However, the specific heat of the BCC lattice with $z=8$ (first neighbors) lies very close to the cubic FCC with $z=12$ and above $z=10$. The FCC lattice with $z=12$ has a C quite close to the HCP with $z=12$ (first and second neighbors). The diamond(16) and EuPdSn2(16) lattices when 16 neighbors are considered also have very similar C. Other exceptions correspond to the FCC(18) case composed of twelve NN and six NNN that has C quite close to the previous mentioned cases of diamond(16) and EuPdSn2(16). Nevertheless, the specific heat is very similar for different structures with the same $z$. Even if the mean-field theory predicts identical behavior for all coordination numbers, without accounting for lattice geometry, this universality seen with QMC across different lattice types is remarkable.
@@
Table 2: Extrapolated critical temperatures as computed with QMC for Hamiltonians in the lattices considered on text. Lattices marked with * have more than third nearest-neighbors (see text). For lattices marked with we report the $N=20$ critical temperature. Uncertainty in $T_{C}$ is around 3%. correspond to a Neel temperature.
IV Effect of anisotropy

### v4

@@ Abstract
We have studied theoretically the specific heat of a large number of non-frustrated magnetic structures described by the Heisenberg model for systems with total angular momentum $J=7/2$, corresponding to the 4f7 configuration of Gd+3 and Eu+2. For a given critical temperature (determined by the magnitude of the exchange interactions), we find that, to a high degree of accuracy, the specific heat is governed by two primary parameters: the effective number of neighbors $z$, which dictates the extent of spatial and quantum fluctuations, and the axial anisotropy KK. The universality of $z$ (its ability to describe specific heat across diverse lattices) holds robustly for systems where exchange interactions do not strongly increase with distance and in the absence of frustration. Otherwise, deviations from universality emerge. Using these two parameters we fit the specific heat of four Gd compounds and two Eu compounds, achieving a remarkable agreement. The present approach enables the extraction of magnetic interaction parameters not accessible through mean-field theory, offering a powerful tool for interpreting specific heat data in 4f7 systems.
@@ I Introduction
However, an exception is expected for compounds containing Gd+3 and Eu+2. Both of these ions correspond to the 4f7 configuration. In this case, the ground-state multiplet, constructed by Hund’s rules, is $^{8}S_{7/2}$. This corresponds to a total spin $S=7/2$, total orbital angular momentum $L=0$ (insensitive to crystal field), and total angular momentum $J=7/2$.
We show in this manuscript that the specific heat of 4f7 systems can be characterized primarily by two parameters: the effective number of neighbors $z$, which accounts for the number of interaction bonds of each ion, and axial anisotropy $K$, which arises due to a small admixture with $L=1$ states.
The thermodynamic behavior of magnetic systems is traditionally analyzed through lattice-specific models, but here we demonstrate that the effective number of neighbors $z$ can provide a universal description of specific heat across diverse magnetic structures. This universality holds robustly for systems where exchange interactions do not strongly increase with distance and in the absence of frustration. In this parameter regime, the specific heat curves of different lattices with the same $z$ collapse onto a single reference curve, enabling a simplified yet accurate interpretation of experimental data. Outside this regime, while the universality of $z$ still provides a qualitative approximation, deviations arise. These deviations highlight the importance of lattice geometry and frustration in systems where the hierarchy of exchange interactions is inverted. Our work builds on these insights to demonstrate that the two-parameter model (defined by $z$ and axial anisotropy $K$) successfully captures the specific heat of 4f7 compounds (Gd3+ and Eu2+) across a wide range of magnetic structures.
@@
In the present study we compute the magnetic specific heat for a wide variety of non‑frustrated (i.e. all exchange bonds can be satisfied simultaneously) lattices using the Heisenberg Hamiltonian with distance‑dependent exchange couplings $H_{0}$ and a single‑ion anisotropy term $H_{A}$:
@@
$\mathbf{S}_{i},S_{iz}$ denote the spin (or total angular momentum) $|S|=7/2$ operators at site $i$ and its projection, $\delta$ labels the non-equivalent space vectors connecting different spins at short distances (nearest and possibly further neighbors), and $J_{\delta}$ is the corresponding exchange interaction. Couplings $J_{\delta}$ are ferromagnetic if negative or antiferromagnetic if positive. We use $S$ instead of $J$ to denote the operators to avoid confusion with the exchange interactions.
The theoretically studied non-frustrated magnetic structures correspond to either ferromagnetic interactions or non-frustrated antiferromagnetic ones. We find that the specific heat of all these structures can be characterized to a large degree of accuracy by two parameters: i) the effective number of neighbors (to be defined more precisely below) and ii) the magnetic anisotropy defined by $H_{A}$. Using these two parameters, and the known value of $T_{C}$, we fit the data for the specific heat of GdNi3Ga9, GdPbBi, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. The corresponding results are presented in Section VII.
The remainder of the paper is organized as follows. In Sec.II.1 we define the effective number of neighbours $z$ and discuss its physical meaning. Sec.II.2 derives the axial anisotropy term for 4f7 ions. Section III describes the quantum Monte Carlo implementation and the finite‑size analysis. In Sec.IV we present the specific‑heat curves for a large set of lattices and demonstrate the universality of $z$ and its limitations in the case of uniform couplings. In Sec. V we show the effect of anisotropy on the specific heat. In Sec. VI we analyze several lattices with non-uniform exchange couplings. SectionVII contains the fits to six experimental compounds, and Sec. VIII summarizes our findings.
II The two main parameters
II.1 The effective number of neighbors
@@
II.2 Axial anisotropy in 4f7 ions
@@
Nevertheless, due to spin-orbit coupling, the ground-state multiplet acquires a small admixture of the $^{6}P_{7/2}$ state with $L=1$. An explicit calculation for Gd+3 in Appendix B of Ref. 18 estimates that the excited multiplet contributes approximately $3\%$ to the ground state. This admixture gives rise to an $L=1$ component in the ground state, which enables its splitting in the presence of crystal fields. This occurs if the symmetry at the rare-earth site is not high. Specifically, if there is only a single axis of higher order, the anisotropy term $H_{A}=C[3L_{z}^{2}-L(L+1)]$ becomes relevant.
@@ III Methods
We have calculated the specific heat of a large number of magnetic structures. The approach used a two-point numerical differentiation to compute the derivative of the internal energy as a function of temperature. The internal energy was calculated using quantum Monte Carlo (QMC) simulations with the default “minbounce” algorithm (implemented in the “dirloop_sse” application in the ALPS libraries [19, 20]). QMC allows us to handle large system sizes and complex lattices while properly accounting for spatial and quantum fluctuations that mean-field approaches miss. Simulations were performed on systems of up to $8\times 30^{3}=216000$ magnetic moments. A typical simulation required approximately 10000 CPU hours per lattice on a standard workstation.
@@
To generate the lattice configurations for the ALPS simulations, we employed the “SUNNY” package [21] in conjunction with crystallographic information files (CIF) obtained from public repositories. SUNNY reads the CIF file and outputs the list of neighbour vectors up to a chosen cutoff. Specifically, CIF files for Fe in simple cubic (SC), body-centered cubic (BCC), hexagonal close-packed (HCP), and face-centered cubic (FCC) structures, as well as C in diamond, MgAgAs (for GdPbBi), Th(CrSi)2 (for GdCu2Ge2), ErNiSi3 (for GdNiSi3), Mg2MnGa3 (for Eu2Pd3Sn) and MgAl2Cu (for EuPdSn2) were retrieved from the Materials Project [22] database. The ErNi3Al9 CIF file was obtained from the Crystallography Open Database [23] using data from Ref. 24. Lattice parameters for GdCu2Ge2 were taken from Ref. [25], for GdPbBi from Ref. [26], for GdNiSi3 from Ref. [27], for Eu2Pd3Sn from Ref. [28], and for EuPdSn2 from Ref. [29]. In each case, the magnetic structure was constructed by assigning a spin moment $S=7/2$ to the respective atoms (Gd or Eu). For all lattices, except HCP, we used a rectangular cuboid supercell geometry as the conventional unit cell, resulting in varying numbers of sites within the supercell, with the diamond lattice requiring the largest (8 sites). Simulations were performed on systems containing $N^{3}$ supercells with periodic boundary conditions, typically with $N=10$, 2020, and $30$. For the critical temperature, results were extrapolated to the thermodynamic limit assuming a $1/N$ dependence.
@@
The well-known sign problem [30] associated with QMC simulations limited our investigation to non-frustrated magnetic structures. At low temperatures, the number of Monte Carlo steps required for statistically reliable results increases exponentially. Consequently, the internal energy was primarily computed in a temperature range between $0.2T_{C}$ and $2T_{C}$, where $T_{C}$ is the critical temperature of the system. In a few cases, simulations were extended to $0.1T_{C}$ at low temperatures. However, the reliability of the results below $0.2T_{C}$ is limited due to finite-size effects. We used $2\times 10^{5}$ Monte Carlo steps for thermal equilibration and $4\times 10^{6}$ steps for measurements. These parameters results in a well converged energy and a smooth specific heat. The resulting error bars (not shown in figures) are close to the line width.
@@
IV Role of the effective number of neighbors
@@
Figure 1: Critical temperature $T_{C}$ in units of the nearest‑neighbour exchange as a function of effective number of neighbors $z$ for several ferromagnetic structures. Finite-size results ($N=10,20,30$) and extrapolated values are shown, with the mean-field value $T_{C}^{\mathrm{MF}}$ indicated by the straight line.
@@
In this section we report our results for the specific heat and critical temperature for several simple lattices as a function of the number of neighbors that interact with a given site. For simplicity, in all lattices considered, all sites are equivalent, and we assume all $J_{\delta}=J_{1}$, where $J_{1}$ is the nearest-neighbor (NN) interaction, up to a certain distance $d$ and 0 for longer distances. Therefore, the coordination number given by Eq. (4) reduces to $\Sigma_{\delta}z_{\delta}. zz$ is given by the total number of neighbors within a sphere of radius $d$. In Section VI we discuss more general cases demonstrating that the results are not severely affected by this assumption.
@@
Most of our studies assume ferromagnetic interactions ($J_{\delta}<0$). Generally, the results for the corresponding non-frustrated antiferromagnetic structures are very similar, particularly for $z>6$, except at very low temperatures, where the specific heat is dominated by magnons and scales as $T^{3/2}$ or $T^{3}$ for ferro- and antiferromagnetic structures respectively (which can be seen in Fig. 6, Sec. VI).
@@
In Fig. 1, we show the critical temperature $T_{C}$ as a function of $z$ for several ferromagnetic structures, including finite-size results and an extrapolated value. Calculations have been done in finite systems containing $N^{3}$ conventional unit cells, with $N=10$, 20, and 30, as explained in Section III. The extrapolated value is also indicated in the figure.
@@
As expected, $T_{C}$ increases almost linearly with $z$ as the mean-field value $T_{C}^{\mathrm{MF}}$ (the effective magnetic field due to the interactions is proportional to $zJ_{1}$, see Section V.1). But it is lower than $T_{C}^{\mathrm{MF}}$, particularly for low $z$ due to the effect of spatial and quantum fluctuations, which is larger for low $z$ (a similar behavior was reported for $S=1/2$ in Ref. [32]).
@@
For large enough $z$, one expects that $T_{C}$ approaches $T_{C}^{\mathrm{MF}}$. However, in finite systems with periodic boundary conditions, the coordination number decreases when the interaction range $d$ becomes comparable to the system size, because neighbors at different distances overlap. This might prevent us from reaching the thermodynamic limit with our method. Arguing as in the introduction that fluctuations decrease as $1/z$ , one should expect that the critical temperature approaches the MF result much faster than the results shown on Fig. 1.
@@
Figure 2: Specific heat $C/T$ versus reduced temperature $T/T_{C}$ for several ferromagnetic structures with different effective number of neighbors $z$. Top panel ($z\leq 12$) and bottom panel ($z\geq 12$) show distinct ranges. The curves for SC, FCC, and HCP structures coincide at $z=18$, while SC, BCC, and EuPdSn2 coincide at $z=26$, explaining the single curves shown for each set. LHc(4) denotes the layered honeycomb lattice with $z=4$. The bottom inset shows finite-size effects for HCP(12) with $N=10$ (black), $N=20$ (red), and $N=30$ (green) in the range $0.5<T/T_{C}<1.5$.
@@
In Fig. 2 we show the results for the specific heat $C$ as a function of temperature for a large number of ferromagnetic structures with different numbers of magnetic neighbors $z$ (the number of nearest neighbors up to third order is shown in Table 1 and its critical temperatures are shown in Table 2). Also shown is the mean-field result which corresponds to the limit $z\rightarrow\infty$. This result shows larger $C$ for $T<T_{C}$, while $C=0$ for $T>T_{C}$. In contrast, the specific heat for a honeycomb layered system with $z=4$ (three nearest neighbors within the plane of the honeycomb layer and one above or below the layer in alternating sites, LHc(4)) shows the smallest specific heat below $T_{C}$ and the largest one above $T_{C}$. Note that the diamond structure, which is distinct from the previous one but also has $z=4$, has a very similar specific heat. The low-temperature shoulder is a Schottky-like effect due to the internal field [33, 14]. The lower panel inset shows that finite size effects are only appreciable near $T_{C}$ as exemplified in the case of the HCP(12) lattice. For this reason, we show specific heat curves from QMC corresponding to $N=20$.
@@
Increasing $z$ from 4 to 6, the specific heat of the simple cubic structure with six NN displays the expected trend of increasing (decreasing) $C$ below (above) $T_{C}$. Increasing $z$ up to 14 using the BCC structure with 6 NN and 8 next-nearest neighbors (NNN), the same trend is observed in general. However, the specific heat of the BCC lattice with $z=8$ (first neighbors) lies very close to the cubic FCC with $z=12$ and above $z=10$. The FCC lattice with $z=12$ has a $C$ quite close to the HCP with $z=12$ (first and second neighbors). The diamond(16) and EuPdSn2(16) lattices when 16 neighbors are considered also have very similar specific heat. Other exception correspond to the FCC(18) case composed of twelve NN and six NNN that has $C$ quite close to the previous mentioned cases of diamond(16) and EuPdSn2(16). Nevertheless, the specific heat is very similar for different structures with the same $z$. Even if the mean-field theory predicts identical behavior for all coordination numbers, without accounting for lattice geometry, this universality seen with QMC across different lattice types with the same $z$ is remarkable.
@@
Table 2: Extrapolated critical temperatures computed with QMC for Hamiltonians in the lattices considered in the text. Lattices marked with * have more than third nearest-neighbors (see text). For lattices marked with we report the $N=20$ critical temperature. AF denotes that antiferromagnetic interactions were used. Uncertainty in $T_{C}$ is below 2%. correspond to a Néel temperature.
V Effect of anisotropy